# Empirical bounds for commuting dilations of free unitaries and the universal commuting dilation constant

### v1

@@ 4.1. Experiment description and collected data
We repeated the experiment described on the left pane of Figure˜4 ten more times. The results are shown in Figure˜5. It seems that all paths are converging to something near $1.44$. Recall that by Theorem˜2.1 and [GS21, Theorem 3.1], we expect to see
@@ 4.2. Code
All numerical experiments were performed using the Julia programming language [BEKS17]. Optimization models were formulated using the Convex.jl package [UMZ+14] and solved with the SCS solver [SCS25]. We used DataFrames.jl [BVK23] and CSV.jl [Q+25] for data handling, and Plots.jl [Plo25] for visualization. Auxiliary scripts for running the series of experiments and visualizing the results were written in Visual Studio Code with the help of GitHub Copilot. The code was run on a MacBook Pro with Apple M4 Pro chip and 24 GB of memory. Our code is available at https://oshalit.net.technion.ac.il/code/.
@@ 4.3. How reliable are the results
The Splitting Conic Solver (SCS) [OCPB16, O’D21] is a well established and tested solver. The optimization finished thousands of runs consistently with status “solved”. It failed in a few cases when we fed the program problems that were too large, in which case no result was recorded.
In order to increase our confidence in the results, we double-checked their reliability by computing the dilation constant for several pairs $(\mathsf{U},\mathsf{N})$ by the algorithm described in Section˜3.2 with both SCS as well as the commercial solver MOSEK [MOS25]. For several tests with $(N,k)=(30,14),(35,14),(40,10)$ and $(45,10)$ we obtained the same dilation constant up to an error of about $0.001$. This is reasonable given SCS’s default tolerance $1\textup{e-}4$ and the fact that we only expect to approximate $c(u_{\mathrm{f}},u_{0})$ to within $0.01$ of its correct value for $k\leq 23$; see Equation˜7.
@@ Appendix A A rigorous bound for the dilation constant $C_{2}(n)$
Denote by $C_{2}(n)$ the dilation constant such that for every pair of contractive $n\times n$ matrices, there exists a dilation to a pair of commuting matrices with norm less than or equal to $C_{2}(n)$. By Ando’s theorem combined with the usual “dilation theory in finite dimensions” yoga [HL21, LS14, MS13] the constant $C_{2}(n)$ is also equal to:
@@
To see this in detail, let $C$ be a constant, and suppose that a pair of $n\times n$ matrices $(A,B)$ has a dilation to $C(S,T)$ where $S,T$ are commuting contractive matrices. By Ando’s theorem [And63], $(S,T)$ has a power dilation to a pair of commuting unitaries $(U,V)$, thus $(A,B)$ has a dilation to $C$ times a pair of commuting unitaries. This shows that $C_{2,u}(n)\leq C_{2}(n)$. Now, if a $(A,B)$ has a dilation to $C$ times a pair of commuting unitaries, then Theorem 1.5 in [HL21] tells us that $(A,B)$ has a dilation to $C$ times a pair $(U^{\prime},V^{\prime})$ of commuting unitary matrices. This shows that $C_{2,u,\operatorname{fin}}(n)\leq C_{2,u}(n)$. On the other hand, it is clear that $C_{2}(n)\leq C_{2,u,\operatorname{fin}}(n)$ and we conclude that

### v2

@@ 4.1. Experiment description and collected data
We repeated the experiment described on the left pane of Figure˜4 ten more times. The results are shown in Figure˜5. It seems that all paths are converging to something near $1.44$. Recall that by Theorem˜2.1 and [19, Theorem 3.1], we expect to see
@@ 4.2. Code
All numerical experiments were performed using the Julia programming language [3]. Optimization models were formulated using the Convex.jl package [39] and solved with the SCS solver [37]. We used DataFrames.jl [8] and CSV.jl [36] for data handling, and Plots.jl [35] for visualization. Auxiliary scripts for running the series of experiments and visualizing the results were written in Visual Studio Code with the help of GitHub Copilot. The code was run on a MacBook Pro with Apple M4 Pro chip and 24 GB of memory. Our code is available at https://oshalit.net.technion.ac.il/code/.
@@ 4.3. How reliable are the results
The Splitting Conic Solver (SCS) [30, 31] is a well established and tested solver. The optimization finished thousands of runs consistently with status “solved”. It failed in a few cases when we fed the program problems that were too large, in which case no result was recorded.
In order to increase our confidence in the results, we double-checked their reliability by computing the dilation constant for several pairs $(\mathsf{U},\mathsf{N})$ by the algorithm described in Section˜3.2 with both SCS as well as the commercial solver MOSEK [28]. For several tests with $(N,k)=(30,14),(35,14),(40,10)$ and $(45,10)$ we obtained the same dilation constant up to an error of about $0.001$. This is reasonable given SCS’s default tolerance $1\textup{e-}4$ and the fact that we only expect to approximate $c(u_{\mathrm{f}},u_{0})$ to within $0.01$ of its correct value for $k\leq 23$; see Equation˜8.
@@ Appendix A A rigorous bound for the dilation constant $C_{2}(n)$
Denote by $C_{2}(n)$ the dilation constant such that for every pair of contractive $n\times n$ matrices, there exists a dilation to a pair of commuting matrices with norm less than or equal to $C_{2}(n)$. By Ando’s theorem combined with the usual “dilation theory in finite dimensions” yoga [22, 26, 27] the constant $C_{2}(n)$ is also equal to:
@@
To see this in detail, let $C$ be a constant, and suppose that a pair of $n\times n$ matrices $(A,B)$ has a dilation to $C(S,T)$ where $S,T$ are commuting contractive matrices. By Ando’s theorem [1], $(S,T)$ has a power dilation to a pair of commuting unitaries $(U,V)$, thus $(A,B)$ has a dilation to $C$ times a pair of commuting unitaries. This shows that $C_{2,u}(n)\leq C_{2}(n)$. Now, if a $(A,B)$ has a dilation to $C$ times a pair of commuting unitaries, then Theorem 1.5 in [22] tells us that $(A,B)$ has a dilation to $C$ times a pair $(U^{\prime},V^{\prime})$ of commuting unitary matrices. This shows that $C_{2,u,\operatorname{fin}}(n)\leq C_{2,u}(n)$. On the other hand, it is clear that $C_{2}(n)\leq C_{2,u,\operatorname{fin}}(n)$ and we conclude that
@@
Acknowledgements. The authors would like to thank the referee for suggesting several corrections and improvements. In particular, Remark 2.8 was stimulated by the referee report.